A 19.5L flask at 15°C contains a mixture of three gases: N2 (2.50mol), He (0.38mol), and Ne (1.34mol). Calculate the partial pressure of neon gas in the mixture.
1
Expert's answer
2020-10-19T14:02:34-0400

Total mole of gas= 4.22 mole

PV = nRT

19.5L×P = 4.22×0.08205746 L atm/k×(15+273)k

Or, P =5.11 atm



Pertial pressure of neon =( 1.34/ total mole)×P

=(1.34/4.22)*5.11

=1.62 atm